# The polynomial set associated with a fixed number of matrix-matrix multiplications

### v1

@@ Abstract
We consider the problem of computing matrix polynomials $p(X)$, where $X$ is a large matrix, with as few matrix-matrix multiplications as possible. More precisely, let $\Pi_{2^{m}}^{*}$ represent the set of polynomials computable with $m$ matrix-matrix multiplications, but with an arbitrary number of matrix additions and scaling operations. We characterize this set through a tabular parameterization. By deriving equivalence transformations of the tabular representation, we establish new methods that can be used to construct elements of $\Pi_{2^{m}}^{*}$ and determine general properties of the set. The transformations allow us to eliminate variables and prove that the dimension is bounded by $m^{2}$. Numerical simulations suggest that this is a sharp bound. Consequently, we have identified a parameterization, which, to our knowledge, is the first minimal parameterization. Furthermore, we conduct a study using computational tools from algebraic geometry to determine the largest degree $d$ such that all polynomials of that degree belong to $\Pi_{2^{m}}^{*}$, or its closure. In many cases, the computational setup is constructive in the sense that it can also be used to determine a specific evaluation scheme for a given polynomial.
@@ 1 Introduction
The computation of matrix functions, in the sense of Higham [13], is a classical problem in numerical linear algebra. We define a matrix function as an extension of a scalar function from $f:\mathbb{C}\rightarrow\mathbb{C}$ to matrices, i.e., $f:\mathbb{C}^{n\times n}\rightarrow\mathbb{C}^{n\times n}$. Important matrix functions like $e^{X}$, $\text{sign}(X)$, and $\sqrt{X}$ are crucial in various fields such as linear ODEs [18], control theory [5], network analysis [7], and quantum chemistry [22]; see [13, Chapter 2] for more applications. Further applications relevant for our setting, appear in problems where the action of a matrix function on a vector $b\in\mathbb{C}^{n}$, i.e., $f(X)b$ is needed, see [9].
@@
In this paper we study methods for computing matrix functions when the input matrix $X\in\mathbb{C}^{n\times n}$ is a very large and dense matrix. In particular, we consider a family of methods that only utilize the following operation types:
  • Linear combination: $Z\leftarrow\alpha X+\beta Y$
  • Multiplication: $Z\leftarrow X\cdot Y$.
@@
A direct consequence of only considering these basic operation types is that this family of methods compute matrix polynomials. Another direct consequence is that the first operation type can be viewed as free in terms of computational cost, since the computational complexity is $O(n^{2})$ and $O(n^{3})$ respectively for the two considered operation types.
@@
We want to study what polynomials can be computed with a given cost. Since, the cost is essentially given by the number of matrix-matrix multiplications, methods with a given fixed cost corresponding to $m$ multiplications, form evaluations of polynomials in the set
@@
where $\Pi_{d}$ is the set of all polynomial of degree $d$ in the usual sense. Here, $\widebar{\Pi}_{d}$ is the closure of $\Pi_{d}$ which is the set of all polynomial of grade $d$, that is, polynomials of degree $\leq d$. In our analysis, we will need to be extra careful concerning the distinction between $\Pi_{k}$ and $\widebar{\Pi}_{k}$. In fact, $\Pi_{2^{m}}^{*}$ is a semi-algebraic set, as we concretize in Section 2.
In an application involving a function $f$, which is not necessarily a polynomial, we often want compute an approximation $p$ of $f$ using a fixed cost. Hence, we want to find the best polynomial approximation in $\Pi_{2^{m}}^{*}$. This is in complete contrast to the classical problem of (scalar) approximation theory where we want to find the best approximation of a given fixed degree. Since $\Pi_{2^{m}}^{*}\neq\widebar{\Pi}_{2^{m}}$, the fixed degree and fixed cost approximation problems are different in character. In order to construct methods for the fixed cost approximation problem we need to understand the set $\Pi_{2^{m}}^{*}$. The objective of this paper is to characterize this set.
Techniques to keep the number of matrix-matrix multiplications low has been studied for decades in the context of matrix functions and matrix polynomials. In Paterson and Stockmeyer’s seminal paper [19] they present an algorithm to compute $p(X)$ for $p\in\widebar{\Pi}_{d}$ in $m=\mathcal{O}(\sqrt{d}$) matrix-matrix multiplications, which was made more precise in [8]. Since they also show that dim$(\Pi_{2^{m}}^{*})=\mathcal{O}(m^{2})$ = $\mathcal{O}(d)$, Paterson and Stockmeyer’s algorithm is optimal in the ordo sense.
@@
The fact that the Paterson and Stockmeyer algorithm is improvable, in particular for moderate $m$, has served as the motivation of a number of recent works. For example, Sastre [24] showed how to compute most polynomials of degree $8$ using only three multiplications, and most polynomials of degree $12$ using only four multiplications (in contrast to the Paterson-Stockmeyer algorithm that handle degrees $6$ and $9$ respectively). The work has been extended in various ways, for example [25], illustrating how an approximation of the Taylor expansion of the exponential can be computed with $m=4$ multiplications; called the approximation 15+ corresponding to a polynomial of degree 16 matching all Taylor coefficients except the last, i.e., the coefficient for $X^{16}$. Corresponding ideas for rational functions were explored in [23]. Further specialziation of the algorithms, in particular for scaling-and-squaring [17, Section 7], has been investigated for the matrix exponential [26], [27], as well as the matrix cosine [28]. Using ideas of efficient polynomial evaluation combined with rational approximations have been used [2, 3], with particular focus on preservation of Lie algebra properties for the matrix exponential. In general, such efficient polynomial methods are (at least in theory) better than the standard implementation of the matrix exponential using a Padé approximant combined with the scaling-and-squaring algorithm [1], although further research is needed to conclusively support this claim.
In the terminology of our framework, considerable parts of the results in the literature discussed above, are examples involving polynomials that are elements of $\Pi_{2^{m}}^{*}$. Despite this research attention, the understanding of this set is far from complete. The basis of our analysis is an evaluation scheme (further explained in Section 2) that gives a parameterization of $\Pi_{2^{m}}^{*}$ with parameters given by a triplet $(A,B,c)$, where $A$ and $B$ are matrices and $c$ a vector. One triplet $(A,B,c)$ corresponds to one polynomial $p\in\Pi_{2^{m}}^{*}$. There are several ways to obtain the same polynomial, i.e., a different triplet $(\hat{A},\hat{B},\hat{c})$ may lead to the same polynomial $p$.
The first set of new results are equivalence transformations. In Section 3, we present procedures to transform a triplet $(A,B,c)$ into another triplet $(\hat{A},\hat{B},\hat{c})$ that yield the same polynomial $p\in\Pi_{2^{m}}^{*}$. Several conclusions can be drawn from the transformations. For example, we prove that the first column of the matrices $A$ and $B$ can be selected as zero without loss of generality. More complex transformations reveal that the $(2,2)$ element of $A$ can also be set to zero. Moreover, we establish a transformation related to the third multiplication, which includes an algebraic condition on the elements of $A$ and $B$. Further analysis shows that the $(3,3)$ element of $A$ can be coupled in simple way to the $(3,3)$ element of $B$ without loss of generality.
@@
Further conclusions from the equivalence transformations are presented in Section 4. Using the transformations we can reduce the number of free parameters that parameterize $\Pi_{2^{m}}^{*}$ and conclude that
@@
This is sharper than bounds presented in, e.g., [19]. Based on a study of the numerical Jacobian, we conjecture that this bound cannot be improved, suggesting that our work present the (to our knowledge) first minimal parameterization of $\Pi_{2^{m}}^{*}$.
@@
Using the transformations and tools from computational algebraic geometry, we solve this problem for $m=4$ and provide conjectures supported by strongly indicative computational results in high precision arithmetic for $m=5$, $m=6$, and $m=7$. Our solutions to (2) are $12$, $20$, $32$, $30$, and $42$ for $m=4$, $m=5$, $m=6$ (complex arithmetic), $m=6$ (real arithmetic), and $m=7$ (complex arithmetic), respectively. The numerical experiments are reproducible and provided in a publicly available repository, including generated code for matlab and Julia.
Although we focus on the determination of dimension and the solution to (2), several of the results seem valuable beyond these specific research questions. The transformation by themselves are constructive, and could be used to improve other properties of the evaluation scheme, e.g., numerical stability. The study of the polynomial subset in Section 5 is constructive. We provide evaluation schemes for the truncated Taylor expansion of the matrix exponential for the given degrees. For example, we compute the degree-$30$ expansion in $m=6$ multiplications, while previous works appear to require $7$ multiplications. Moreover, the software that we provide is able to compute the evaluation scheme coefficients for many given polynomials.
@@ 2.1 Definition of the evaluation scheme
Fundamental to the results of this paper is an evaluation scheme to construct elements of $\Pi_{2^{m}}^{*}$. Such evaluation schemes have been presented in various forms in [19, Section 2], [2, section 3], and also [14], where in the latter they are referred to as degree-optimal polynomials. The scheme involves parameters stored in matrices $A$ and $B$, and vector $c$. Matrices $A$ and $B$ contain coefficients for linear combinations, where each row $k$ corresponds to the linear combinations associated with the $k$th multiplication. As we perform $m$ multiplications, these matrices have $m$ rows. The vector $c$ contains coefficients for linear combinations that are used after the multiplications have been completed.
In particular, consider the triplet $(A,B,c)\in\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m+2}$ and associate the following sequence of operations involving precisely $m$ matrix-matrix multiplications. Define $Q_{1}=I$ and $Q_{2}=X$. Then, iterate the process to generate $Q_{3},\ldots,Q_{m+2}$ using the elements of matrices $A$ and $B$ as follows:
@@
Furthermore, compute the output of the scheme by forming the linear combination corresponding to the $c$-vector:
@@
Thus, a given triple $(A,B,c)$, with $A$ and $B$ being Hessenberg matrices, defines a polynomial $p\in\Pi_{2^{m}}^{*}$.
Note that all evaluation schemes can be expressed in this manner, given that for each multiplication step, we use all preceding information accessible at no additional computational expense. For instance, in the second multiplication step, we determine $Q_{4}$ by combining linearly the prior multiplications $Q_{3}$ along with the inputs $Q_{2}$ and $Q_{1}$. Hence, a triplet $(A,B,c)$ parameterizes $\Pi_{2^{m}}^{*}$ through the expressions in (3) and (4). Detailed examples illustrating the representation of monomial evaluation or Horner evaluation can be found in [14].
@@
To see how the tables represent standard evaluation schemes, such as monomial evaulation, Horner evaluation, and Paterson-Stockmeyer evaluation, see [14, Section 3]. In the following, we show how to evaluate $X^{7}+\epsilon X^{8}$ in three multiplications. Consider the coefficient triplet:
(8) $[A|B]$ $=\left[\begin{array}[]{c c c c | c c c c}0&1&&&0&1&&\\ 0&\frac{1}{2\epsilon}&1&&0&0&1&\\ 0&\frac{7+8\epsilon^{2}-16\epsilon^{4}}{128\epsilon^{5}}&-\frac{1}{8\epsilon^{2}}-\frac{1}{2}&1&0&\frac{-7+8\epsilon^{2}+16\epsilon^{4}}{128\epsilon^{5}}&-\frac{1}{8\epsilon^{2}}+\frac{1}{2}&1\end{array}\right]$
(10) $c$ $=\left[\begin{array}[]{ccccc}0&0&\frac{49-288\epsilon^{4}+256\epsilon^{8}}{16384\epsilon^{9}}&\frac{-5+16\epsilon^{4}}{64\epsilon^{3}}&\epsilon\end{array}\right].$
For this triplet $(A,B,c)$, we obtain the polynomial $p(X)=X^{7}+\epsilon X^{8}$ using the evaluations specified in (3)-(4). We observe that $X^{7}$ is as a limit point of $\Pi_{2^{3}}^{*}$ corresponding to $\epsilon\rightarrow 0$, although $X^{7}$ requires four multiplications. This example illustrates the complexity of $\Pi_{2^{m}}^{*}$ for $m=3$. For higher $m$, this complexity becomes even more intricate.
@@ 2.2 The semi-algebraic set perspective of $\Pi_{2^{m}}^{*}$
For our study, we need concepts from algebraic geometry. Let $\mathbb{K}$ be the field of the parameters in the evaluation, i.e., in most applications $\mathbb{K}=\mathbb{C}$ or $\mathbb{K}=\mathbb{R}$. Let $\mathbb{X}$ be the product space associated with the parameters. That is $\mathbb{X}:=\mathbb{K}^{m(m+3)/2}\times\mathbb{K}^{m(m+3)/2}\times\mathbb{K}^{m+2}=\mathbb{K}^{s}$, where $s=m^{2}+4m+2$ is the number of parameters in the evaluation.
If we denote the evaluation scheme (3)-(4) by the map $\Phi:(A,B,c)\mapsto p\in\widebar{\Pi}_{d}$, we can describe $\Pi_{2^{m}}^{*}$ as its image
@@
By construction, $\Phi$ depends polynomially on the parameters. This can be concluded from the property that the image of a smooth map, with a vector space as its domain, is inherently semi-algebraic.
The set is indeed semi-algebraic and not necessarily an algebraic variety. Let $\widebar{\cdot}$ denote topological closure in a Zariski sense. Then, in relation to the example (8) above, we see that $\epsilon=0$ is not a valid choice, and indeed $X^{7}\not\in\Pi_{2^{3}}^{*}$. However, since $\epsilon=0$ is a limit point, we have $X^{7}\in\widebar{\Pi_{2^{3}}^{*}}=\widebar{\Pi}_{2^{3}}$; at least for $\mathbb{K}=\mathbb{C}$ due to [32, Corollary 4.20].
If $\varphi_{0},\ldots,\varphi_{d}$ denote a basis of the ambient space $\mathbb{K}[x]_{d}$, e.g., as in our simulations a monomial basis, and let $J$ be the Jacobian of the map $\Phi$ with an output expressed in this basis. Following the standard definitions in algebraic geometry [29], the dimension of a semi-algebraic set is given by the rank of $J$ at any non-singular point. Most points in the image of a smooth map are non-singular; therefore for essentially all $(A,B,c)$ we have
@@
where $J\in\mathbb{K}^{(d+1)\times s}$ and $d=2^{m}$ and $s$ is the number of parameters in the method.
@@
(12) $\dim\Pi_{2^{m}}^{*}\leq s=m^{2}+4m+2$
In this paper we improve this bound by reducing the number of free parameters and in Section 4 conclude that $m^{2}$ seems to be the exact dimension, for $m>3$.
Here we note that the dimension may be different for $\mathbb{K}=\mathbb{C}$ or $\mathbb{K}=\mathbb{R}$. Most of the results in this paper are valid for either choice, but in situations where we explicitly refer to the Zariski closure, we assume $\mathbb{K}=\mathbb{C}$ unless otherwise stated. In the simulations in Section 5 we focus on $\mathbb{K}=\mathbb{C}$ with some additional observations for $\mathbb{K}=\mathbb{R}$.

### v3

@@ Abstract
We consider the problem of computing matrix polynomials $p(X)$, where $X$ is a large dense matrix, with as few matrix-matrix multiplications as possible. More precisely, let $\Pi_{2^{m}}^{*}$ represent the set of polynomials computable with $m$ matrix-matrix multiplications, but with an arbitrary number of matrix additions and scaling operations. We characterize this set through a tabular parameterization. By deriving equivalence transformations of the tabular representation, we establish new methods that can be used to construct elements of $\Pi_{2^{m}}^{*}$ and determine general properties of the set. The transformations allow us to eliminate variables and prove that the dimension is bounded by $m^{2}$, which is subsequently proven to be sharp, i.e., $\dim(\Pi_{2^{m}}^{*})=m^{2}$. Consequently, we have identified a parameterization that, to the best of our knowledge, is the first minimal parameterization. We also conduct a study using computational tools from algebraic geometry to determine the largest degree $d$ such that all polynomials of that degree belong to $\Pi_{2^{m}}^{*}$, or its closure. In many cases, the computational setup is constructive in the sense that it can also be used to determine a specific evaluation scheme for a given polynomial.
@@ 1 Introduction
The application that motivates the research question in this paper, is the computation of matrix functions in the sense of Higham [13], which is a classical problem in numerical linear algebra. We define a matrix function as an extension of a scalar function from $f:\mathbb{C}\rightarrow\mathbb{C}$ to matrices, i.e., $f:\mathbb{C}^{n\times n}\rightarrow\mathbb{C}^{n\times n}$. Important matrix functions like $e^{X}$, $\text{sign}(X)$, and $\sqrt{X}$ are crucial in various contexts such as linear ODEs [18], control theory [5], network analysis [7], and quantum chemistry [22]; see [13, Chapter 2] for more applications. Further applications relevant for our setting appear in problems where the action of a matrix function on a vector $b\in\mathbb{C}^{n}$, i.e., $f(X)b$ is needed, see [9].
@@
In this paper we study methods for computing matrix functions when the input matrix $X\in\mathbb{C}^{n\times n}$ is very large and dense. In particular, we consider a family of methods that only utilize the following two operation types:
  • Linear combination of two matrices $Z\leftarrow\alpha X+\beta Y$
  • Multiplication of two matrices $Z\leftarrow X\cdot Y$.
@@
A direct consequence of considering only these basic operation types is that this family of methods computes matrix polynomials. Another direct consequence is that the first operation type can be viewed as free in terms of computational cost, since the computational complexity is $O(n^{2})$ and $O(n^{3})$ respectively for the two considered operation types.
@@
We want to study the polynomials that can be computed with a given cost. Since the cost is essentially given by the number of matrix-matrix multiplications, methods with a given fixed cost corresponding to $m$ multiplications form evaluations of polynomials in the set
(1) $\Pi_{2^{m}}^{*}:=\{p\in\widebar{\Pi}_{2^{m}}\hskip-3.0pt:\hskip-1.0pt\text{$p(X)$ is computable with $m$ matrix-matrix multiplications}\},$
where $\Pi_{d}$ is the set of all polynomials of degree $d$ in the usual sense. Here, $\widebar{\Pi}_{d}$ is the closure of $\Pi_{d}$, that is, the set of polynomials of degree $\leq d$.
Although our application stems from matrix polynomials, the set $\Pi_{2^{m}}^{*}$ is, in fact, a univariate polynomial set, not a matrix polynomial set; it is a univariate semi-algebraic set, as we concretize in Section 2. The set can be defined more abstractly as follows. Let $\Pi=\mathbb{K}[x]$ be the polynomial ring over a field $\mathbb{K}$; typically the polynomial coefficients are scalars with $\mathbb{K}=\mathbb{C}$ or $\mathbb{K}=\mathbb{R}$, such that $\Pi_{d}\subset\Pi$ is the vector space of polynomials in $\mathbb{K}[x]$ of degree at most $d$. We make the following assumptions about computations involving elements of $\Pi$:
  • Linear combination of two polynomials $r\leftarrow\alpha p+\beta q$ is considered computationally free, where $p,q\in\Pi$ are polynomials, and $\alpha,\beta\in\mathbb{K}$ are scalars.
  • Multiplication of two polynomials $r\leftarrow p\cdot q$, where $p,q\in\Pi$, incurs unit computational cost, independent of the degrees of $p$ and $q$.
Now, if we fix a computational budget to $m$, i.e., fix the number of non-scalar multiplications, the total space of computable polynomials is restricted. Since each multiplication can at most double the degree (i.e., $\deg(fg)\leq\deg(f)+\deg(g)$), the maximal degree reachable using $m$ multiplications is bounded by $2^{m}$. The definition of $\Pi^{*}_{2^{m}}\subset\Pi_{2^{m}}$ in (1) corresponds to all polynomials in $\Pi$ that can be computed using at most $m$ non-scalar multiplications, and an arbitrary number of free linear combination operations. Although the matrix polynomial evaluation application is our main source of interest, there are other applications, for example, when the matrix $X$ is replaced by a (scalar) polynomial; see the discussion and references in Section 6.
We also wish to stress the difference of this setting in comparison to the approximation theory assumptions. A common application of matrix polynomials involves the approximation of a non-polynomial function $f$. In such applications, one often wants to compute a polynomial approximation $p$ of $f$ using a fixed cost. Hence, we want to find the best approximation in $\Pi_{2^{m}}^{*}$. This is in complete contrast to the classical problem in (scalar) approximation theory, where one seeks the best approximation of a given fixed degree. Since $\Pi_{2^{m}}^{*}\neq\widebar{\Pi}_{2^{m}}$ in general, the fixed degree and fixed cost approximation problems are different in character. In order to construct methods for the fixed cost approximation problem we need to understand the set $\Pi_{2^{m}}^{*}$. The objective of this paper is to characterize this set.
@@
The fact that the Paterson and Stockmeyer algorithm is often improvable has served as the motivation for a number of recent works. For example, Sastre [24] showed how to improve the Paterson–Stockmeyer method in general; the result shows how to compute most polynomials of degree $8$ using only three multiplications, and most polynomials of degree $12$ using only four multiplications (in contrast to the Paterson–Stockmeyer algorithm that handles degrees $6$ and $9$, respectively). The work has been expanded in various ways (e.g., [25]). This research demonstrates how an approximation of a general function, such as the Taylor expansion of the exponential, can be computed using $m=4$ multiplications. Specifically, the approximation known as 15+ corresponds to a polynomial of degree 16 that matches all Taylor coefficients except for the last, i.e., the coefficient for $X^{16}$. Related concepts for rational functions were examined in [23]. Further refinements of the algorithms, especially regarding the scaling-and-squaring method [18, Section 7], have been investigated for both the matrix exponential [27] and the matrix cosine [29] as well as in [28]. Ideas for efficient polynomial evaluation combined with rational approximations have been used in [2, 3], with particular focus on preservation of Lie algebra properties for the matrix exponential. In general, these types of efficient polynomial methods are (at least in theory) better than the standard implementation of the matrix exponential using a Padé approximant combined with the scaling-and-squaring algorithm [1], although further research is needed to definitively support this claim.
[…]
@@
The starting point of Section 4 are conclusions of the equivalence transformations. Using the transformations we can reduce the number of free parameters that parameterize $\Pi_{2^{m}}^{*}$ and conclude that
@@
This bound is sharper than those presented in, for example, [19]. Moreover, we prove that this is optimal and, to our knowledge, our parameterization therefore forms the first minimal parameterization of unreduced evaluation schemes in $\Pi_{2^{m}}^{*}$.
@@
Using the transformations and tools from computational algebraic geometry, we solve this problem for $m=4$ and provide conjectures supported by strongly indicative computational results in high-precision arithmetic for $m=5$, $m=6$, and $m=7$. Our solutions to (2) are $12$, 2020, $32$, $30$, and $42$ for $m=4$, $m=5$, $m=6$ (complex arithmetic), $m=6$ (real arithmetic), and $m=7$ (complex arithmetic), respectively. The numerical experiments are reproducible and provided in a publicly available repository, including generated code for Matlab and Julia.
While the simulations focus on the determination of the dimension of $\Pi_{2^{m}}^{*}$ and the solution to (2), other aspects of the findings in this paper may be significant beyond these specific research questions. The transformations themselves are constructive and may be used to improve various properties of the evaluation scheme, such as numerical stability. The study of the polynomial subset in Section 5 is also constructive. We provide evaluation schemes for the truncated Taylor expansion of the matrix exponential at specified degrees. For instance, we achieve the degree-$30$ expansion using $m=6$ multiplications, whereas prior works require $7$ multiplications [27]. Moreover, our simulations are applicable to several different functions, and the software we provide can be used to compute the evaluation scheme coefficients for polynomials other than those reported in this paper.
@@ 2.1 Definition of the evaluation scheme
Fundamental to the results of this paper is an evaluation scheme for constructing elements of $\Pi_{2^{m}}^{*}$. Such evaluation schemes have been presented in various forms in [19, Section 2], [2, section 3], and [14], where in [14] they are referred to as degree-optimal polynomials. The evaluation scheme involves parameters stored in the matrices $A$ and $B$, and the vector $c$. Matrices $A$ and $B$ contain coefficients for linear combinations, where each row $k$ corresponds to the linear combinations associated with the $k$th multiplication. As we perform $m$ multiplications, these matrices have $m$ rows. The vector $c$ contains coefficients for linear combinations that are used after the multiplications have been completed.
In particular, consider the triplet $(A,B,c)\in\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m+2}$ and associate the following sequence of operations involving exactly $m$ matrix-matrix multiplications. Define $Q_{1}=I$ and $Q_{2}=X$. Then, we iterate the process to generate $Q_{3},\ldots,Q_{m+2}$ using the elements of matrices $A$ and $B$ as follows:
@@
Furthermore, we compute the output of the scheme by forming the linear combination corresponding to the $c$-vector:
@@
Thus, a given triple $(A,B,c)$, with $A$ and $B$ being matrices, defines a polynomial $p\in\Pi_{2^{m}}^{*}$.
Note that all evaluation schemes can be expressed in this manner, due to the fact that for each multiplication step we use all preceding information available. For instance, in the second multiplication step we determine $Q_{4}$ by computing the product of two linear combinations of $Q_{3}$, $Q_{2}$ and $Q_{1}$. Hence, a triplet $(A,B,c)$ parameterizes $\Pi_{2^{m}}^{*}$ through the expressions in (3) and (4).
@@
To see how the tables represent standard evaluation schemes, such as monomial evaluation, Horner evaluation, and Paterson–Stockmeyer evaluation, see [14, Section 3]. In the following, we show how to evaluate
$p(X)=X^{7}+\epsilon X^{8},$
in three multiplications. Consider the coefficient triplet:
(5d) $[A|B]$ $=\left[\begin{array}[]{c c c c | c c c c}0&1&&&0&1&&\\ 0&\frac{1}{2\epsilon}&1&&0&0&1&\\ 0&\frac{7+8\epsilon^{2}-16\epsilon^{4}}{128\epsilon^{5}}&-\frac{1}{8\epsilon^{2}}-\frac{1}{2}&1&0&\frac{-7+8\epsilon^{2}+16\epsilon^{4}}{128\epsilon^{5}}&-\frac{1}{8\epsilon^{2}}+\frac{1}{2}&1\end{array}\right]$
(5f) $c$ $=\left[\begin{array}[]{ccccc}0&0&\frac{49-288\epsilon^{4}+256\epsilon^{8}}{16384\epsilon^{9}}&\frac{-5+16\epsilon^{4}}{64\epsilon^{3}}&\epsilon\end{array}\right].$
For this triplet $(A,B,c)$, we obtain the polynomial $p(X)=X^{7}+\epsilon X^{8}$ using the evaluations specified in (3)-(4). We observe that $X^{7}$ is a limit point of $\Pi_{2^{3}}^{*}$ corresponding to $\epsilon\rightarrow 0$, even though $X^{7}$ requires four multiplications. This example illustrates the complexity of $\Pi_{2^{m}}^{*}$ for $m=3$. As $m$ increases, the complexity becomes even more intricate.
@@ 2.2 The semi-algebraic set perspective of $\Pi_{2^{m}}^{*}$
For our study, we need concepts from algebraic geometry. Let $\mathbb{K}$ be the field of the parameters in the evaluation. Let $\mathbb{X}$ be the product space associated with the parameters. That is, $\mathbb{X}:=\mathbb{K}^{m(m+3)/2}\times\mathbb{K}^{m(m+3)/2}\times\mathbb{K}^{m+2}=\mathbb{K}^{s}$, where $s=m^{2}+4m+2$ is the number of parameters in the evaluation.
If we denote the evaluation scheme (3)–(4) by the map $\Phi:(A,B,c)\mapsto p\in\widebar{\Pi}_{d}$, we can describe $\Pi_{2^{m}}^{*}$ as its image
@@
By construction, $\Phi$ is algebraic, since it depends polynomially on the parameters. It follows from the Tarski–Seidenberg theorem [17, Theorem 8.6.6] that $\Pi_{2^{m}}^{*}$ is a semi-algebraic set, since it is the image of an algebraic map over a vector space.
@@
If $\varphi_{0},\ldots,\varphi_{d}$ denotes a basis of the ambient space $\mathbb{K}[x]_{d}$, e.g., as in our simulations a monomial basis, and let $J$ be the Jacobian of the map $\Phi$ with an output expressed in this basis. Following the standard definitions in algebraic geometry [30], the dimension of a semi-algebraic set is given by the rank of $J$ at any non-singular point. Most points in the image of a smooth map are non-singular; therefore for almost all $(A,B,c)$ we have
$\dim\Pi_{2^{m}}^{*}=\mathop{\operator@font rank}\nolimits J,$
where $J\in\mathbb{K}^{(d+1)\times s}$, $d=2^{m}$ and $s$ is the number of parameters in the method.
@@
(7) $\dim\Pi_{2^{m}}^{*}\leq s=m^{2}+4m+2.$
In this paper we improve this bound by reducing the number of free parameters and in Section 4 we subsequently conclude that $m^{2}$ is the exact dimension, for $m>3$.
We note that the properties of the set $\Pi_{2^{m}}^{*}$ differ from an algebraic perspective when considering $\mathbb{K}=\mathbb{C}$ versus $\mathbb{K}=\mathbb{R}$. Most results in this paper are applicable to both cases; however, when explicitly discussing the Zariski closure, we will assume $\mathbb{K}=\mathbb{C}$ unless stated otherwise. In the simulations presented in Section 5, we concentrate on $\mathbb{K}=\mathbb{C}$ and provide additional observations for $\mathbb{K}=\mathbb{R}$.